A soft drink company holds a contest in which a prize may be revealed on the inside of the bottle cap. The probability that each bottle cap reveals a prize is 0.2 and winning is independent from one bottle to the next. What is the probability that a customer must open three or more bottles before winning a prize
(0.2)(0.2)(0.8) = 0.032
(0.8)(0.8)(0.2) = 0.128
(0.8)(0.8) = 0.64
1- (0.2)(0.2)(0.8) = .968
Thanks,
Kim

Answer by stanbon(75887)

A soft drink company holds a contest in which a prize may be revealed on the inside of the bottle cap. The probability that each bottle cap reveals a prize is 0.2 and winning is independent from one bottle to the next.
---------------------------------------------------------
What is the probability that a customer must open three or more bottles before winning a prize
---
The customer cannot win on the 1st or 2nd bottle.

(0.2)(0.2)(0.8) = 0.032
(0.8)(0.8)(0.2) = 0.128
(0.8)(0.8) = 0.64
1- (0.2)(0.2)(0.8) = .968
----
P(wins on the 3rd bottle) = 0.8*0.8*0.2 = 0.128
===================================================
Cheers,
Stan H.